1. Solve the spherical triangle ABC right angled C given c= 56°13´ and b=48°30´.


1
Expert's answer
2021-11-30T11:21:50-0500

"S=\frac{1}{2}absinC"

"a+b+c=180^o"

"a=180-c-b"

"a=180-56^o13"'"-48^o30"'

"a=71^o32"'